OpenStudy (anonymous):

Given right triangle ABC . To the nearest degree, find the measure of angle B. 43° 47° 70°

OpenStudy (alexandervonhumboldt2):

sin(b)=4.8/5.1 b=sin-1(4.8/5.1)

OpenStudy (anonymous):

thnx